OpenStudy (anonymous):

0.02(y-8)+0.10y=0.04-0.03(30)

OpenStudy (anonymous):

0.02y-0.16+0.10y=0.04-0.09 0.12y-0.16=0.04-0.09 0.12y=0.04-0.09+0.16 0.12y=0.21 y=(0.21)/(0.12) =1.75

OpenStudy (anonymous):

There is a typo, it should be like this: 0.02y-0.16+0.10y=0.04-0.9 0.12y-0.16=0.04-0.9 0.12y=0.04-0.9+0.16 0.12y=0.25-0.9 0.12y=-0.65 y=(-0.65)/(0.12) =-5.41666666...